OpenStudy (anonymous):

if you have a system which has a solution set x=9t-11 y=2-5t z=t where t is any real number, is the solution set a subspace of R3. I am thinking yes as you can represent this as (x,y,z)=(-11,2,0)+t(9,-5,1) where the point and the vector(9,-5,1) represents a subspace of points in R3? Is my thought process right here?

OpenStudy (anonymous):

any thoughts?

OpenStudy (dan815):



OpenStudy (anonymous):

my thoughts are correcT?

OpenStudy (dan815):

ya cuz it can be represented in that parametric equation

OpenStudy (anonymous):

i wasn't quite sure that a point can also be considered as a subspace along with a vector

OpenStudy (dan815):

ok i dunno i just wanted to try and answer lol

OpenStudy (dan815):

i dunno maybe?? what if u say this is R4, with x , y , z and t?

OpenStudy (anonymous):

i was thinking that the point doesn't have a t value for that corresponding vector. therefore if you consider that point and the vector then it would make every point in r3? not sure

OpenStudy (dan815):

id say R3 because, u can represent parametric equations in 3 space even if its x=x(t),y=y(t),z=z(t)

OpenStudy (anonymous):

mmk. cheers bro

OpenStudy (dan815):

ok :D

OpenStudy (amistre64):

a subspace has certain requirements

OpenStudy (amistre64):

is the zero vector contained in the equation?

OpenStudy (anonymous):

um nope

OpenStudy (amistre64):

if there is no zero vector, then it is not a subspace ...

OpenStudy (anonymous):

oh stupid me

OpenStudy (dan815):

what is a subspace lol

OpenStudy (dan815):

what course is that?

OpenStudy (amistre64):

its usually a part of linear algebra

OpenStudy (dan815):

oh ok i learn that next year i think or the year after

OpenStudy (dan815):

we did learn about some matrices stuff, thats about it

OpenStudy (anonymous):

yea i tots forgot that! cheers though for the link!

OpenStudy (amistre64):

youll do fine :) good luck

OpenStudy (anonymous):

got exam on 22nd of june. get pumped

OpenStudy (dan815):

are u doing summer school

OpenStudy (anonymous):

nah i'm in australia

OpenStudy (dan815):

o cool :)

OpenStudy (anonymous):

yea first year uni

OpenStudy (dan815):

we never learned this in first year uni.. i wonder whats wrong with my uni lol

OpenStudy (dan815):

all we did was learn about eigen values and stop there, and then how to solve odes with that stuff

OpenStudy (anonymous):

ahah yea i heard that in america you don't even do complex numbers in final year of high school.well thats how the myth goes in australia..aha

OpenStudy (dan815):

lol

OpenStudy (dan815):

we did see it in gr 11 and 12 i think, but very light

OpenStudy (anonymous):

sweet

OpenStudy (dan815):

what kinda other math stuff have u done more than us O_O

OpenStudy (dan815):

im in Canada though so i dunno about USA

OpenStudy (dan815):

have you done multivariable calculus yet?

OpenStudy (anonymous):

not sure what that is. we've done system of differential equations, slope fields etc loads of vectors

OpenStudy (anonymous):

yea eigen vectors

OpenStudy (dan815):

oh ok i think u guys are focusing more on matrices stuff

OpenStudy (anonymous):

we do alot of trig calc and integration

OpenStudy (dan815):

we focused a bit more on integration stuff. like vector calculus and stuff, but we got a linear algebra course that everyones gotta take in 2nd or 3rd year and thats gonna cover all that matrices stuff at once

OpenStudy (anonymous):

ah right. our maths is divided up into linear algebra and calculus 2 lessons of algebra and 2 lessons of calc a week